Grade 11General Physics

if a velocity of light(c), acceleration due to gravity(g) and pressure (p) are taken as the fundamental quantities , then the dimension of gravitational constant is

Vandana Soan
10 Years agoGrade 11

1 Answer

Jigar Marvaniya
ApprovedApproved Tutor Answer10 Years ago
Gravitational constant = M^-1L^3T^-2
So, M-1​L3​T-2​ = [LT-1​]a​ [LT-​2​]b​ [ML-1​T-2​]c​
M-1​L3​T-2​ =  Mc​ La+b-c​ T-1a-2b-2c​  
By comparing
c=-1, b=2,a=0
Gravitational constant= c^0 g^2 p^-1